# GazeGaussian: High-Fidelity Gaze Redirection with 3D Gaussian Splatting

### v1

@@ Abstract
Gaze estimation encounters generalization challenges when dealing with out-of-distribution data. To address this problem, recent methods use neural radiance fields (NeRF) to generate augmented data. However, existing methods based on NeRF are computationally expensive and lack facial details. 3D Gaussian Splatting (3DGS) has become the prevailing representation of neural fields. While 3DGS has been extensively examined in head avatars, it faces challenges with accurate gaze control and generalization across different subjects. In this work, we propose GazeGaussian, a high-fidelity gaze redirection method that uses a two-stream 3DGS model to represent the face and eye regions separately. By leveraging the unstructured nature of 3DGS, we develop a novel eye representation for rigid eye rotation based on the target gaze direction. To enhance synthesis generalization across various subjects, we integrate an expression-conditional module to guide the neural renderer. Comprehensive experiments show that GazeGaussian outperforms existing methods in rendering speed, gaze redirection accuracy, and facial synthesis across multiple datasets. We also demonstrate that existing gaze estimation methods can leverage GazeGaussian to improve their generalization performance. The code will be available at: https://ucwxb.github.io/GazeGaussian/.
@@ Introduction
Figure 1: GazeGaussian for gaze redirection: Given an input image, GazeGaussian deforms face and eye Gaussians from canonical space to generate high-fidelity head images with accurate gaze redirection.
@@
Gaze estimation is a fundamental component across various applications (Andrist et al. 2014; Padmanaban et al. 2017; Mavely et al. 2017), yet current estimators (Cheng et al. 2024; Cheng and Lu 2022; Xu, Wang, and Lu 2023) often struggle to generalize effectively to out-of-distribution data. To address this, recent approaches (Ruzzi et al. 2023; Yin et al. 2024; Wang et al. 2023) have started exploring gaze redirection, which manipulates the gaze in an input image toward a target direction. This process generates augmented data to enhance the generalization capabilities of gaze estimators.
@@
Earlier methods (Ganin et al. 2016; Yu and Odobez 2020; Zhang, Liu, and Lu 2022; Zhang, Sugano, and Bulling 2018) formulate gaze redirection as a 2D image manipulation task, relying on deep learning techniques to warp eye regions of the image toward the target gaze direction. However, these 2D approaches overlook the inherently 3D nature of head and gaze manipulation, often resulting in poor spatial consistency and limited synthesis fidelity. With advancements in Neural Radiance Fields (NeRF) (Mildenhall et al. 2020) and its variants (Wang et al. 2021b; Wei et al. 2024), several methods (Zielonka, Bolkart, and Thies 2023; Grassal et al. 2022; Zheng et al. 2023; Hong et al. 2022a) have achieved 3D dynamic head representation and high-fidelity avatar synthesis. Meanwhile, to enable precise control of gaze direction, recent research (Ruzzi et al. 2023; Yin et al. 2024; Wang et al. 2023) has introduced approaches that decouple the face and eye regions, modeling each with separate neural fields to achieve accurate gaze redirection.
@@
As NeRF-based methods are hindered by high computational demands, 3D Gaussian Splatting (Kerbl et al. 2023) and its variants (Lu et al. 2024; Huang et al. 2024; Wang et al. 2024) achieve impressive rendering quality with significantly faster training speeds. Recent research (Xiang et al. 2024; Qian et al. 2024; Xu et al. 2024) has applied these methods to 3D head animation, typically using face-tracking (Tran and Liu 2018; Zielonka, Bolkart, and Thies 2022) parameters to model dynamic 3D head representations. However, existing 3DGS-based approaches neglect the accurate control of gaze direction and struggle to generalize across different subjects, limiting their effectiveness for gaze redirection.
@@
To address the above issues, we propose GazeGaussian, a high-fidelity gaze redirection method that leverages a two-stream 3D Gaussian Splatting (3DGS) model to represent the face and eye regions, respectively. To the best of our knowledge, this is the first integration of 3DGS into gaze redirection tasks. An overview is shown in Fig. 1.
@@
GazeGaussian begins by initializing the two-stream 3DGS model using a pre-trained neutral mesh on the training dataset. This mesh is divided into distinct regions for the face and eyes. By employing gaze direction and face tracking codes, we optimize a deformation field for the face and a rotation field for the eyes, allowing us to adjust the neutral Gaussians accordingly. To achieve precise eye rotation aligned with the target gaze, we present a novel Gaussian Eye Rotation Representation (GERR). In contrast to methods like GazeNeRF that implicitly alter feature maps, GazeGaussian explicitly adjusts the position of Gaussians in the eye branch according to the desired gaze direction, utilizing the controllable nature of 3DGS. To address possible errors in gaze direction, GazeGaussian develops an eye rotation field to enhance redirection accuracy. The two-stream Gaussians are rasterized into high-level features and sent to the neural renderer. Finally, to enhance synthesis generalization across different subjects and preserve facial details, we employ an expression-guided neural renderer (EGNR) to synthesize the final gaze-redirection images.
@@
  • We introduce GazeGaussian, the first 3DGS-based gaze redirection pipeline, achieving precise gaze manipulation and high-fidelity head avatar synthesis.
@@
  • To enhance the synthesis generalization of 3DGS, we design an expression-guided neural renderer (EGNR) to retain facial details across various subjects.
@@
Gaze Redirection. Gaze redirection is the task of manipulating the gaze direction of a face image to a target direction while preserving the subject’s identity and other facial details. Earlier approaches for gaze redirection include novel view synthesis (Criminisi et al. 2003; Kuster et al. 2012; Giger et al. 2014), eye-replacement (Qin et al. 2015; Shu et al. 2016), and warping-based methods (Ganin et al. 2016; Kononenko and Lempitsky 2015; Wood et al. 2017). However, these methods are limited by person-specific data requirements, restricted redirection range, and artifact introduction. To further improve gaze redirection, recent studies (Park et al. 2019; Zheng et al. 2020; He et al. 2019; Xia et al. 2020) have employed neural network-based generative models. STED (Zheng et al. 2020), building on the FAZE (Park et al. 2019), introduces a self-transforming encoder-decoder that generates full-face images with high-fidelity control over gaze direction and head pose. Effective gaze redirection should account for both the 3D nature of eyeball rotation and the deformation of surrounding facial regions. With advancements in Neural Radiance Fields (NeRF) (Mildenhall et al. 2020), several studies (Li et al. 2022; Ruzzi et al. 2023; Yin et al. 2024; Wang et al. 2023) have aimed to model the complex rotation of the eyeball. GazeNeRF (Ruzzi et al. 2023) employs a two-stream MLP architecture to separately model the face only and eye regions, achieving improved gaze redirection performance.
@@
However, these methods are hindered by substantial computational demands and limited rendering efficiency. Additionally, gaze manipulation occurs at the feature map level and remains an implicit approach. In contrast, GazeGaussian allows for explicit control over eye rotations, improving gaze redirection accuracy and accelerating the synthesis process.
@@
Head Avatar Synthesis. The synthesis of head avatars has garnered considerable attention in recent years. FLAME (Li et al. 2017) is a parameterized 3D head model that maps parameters of shape, expression, and pose onto 3D facial geometry, allowing for realistic and controllable head avatar generation. Many subsequent works (Cudeiro et al. 2019; Ranjan et al. 2018; Fan et al. 2022; Peng et al. 2023a, b; Chen et al. 2023) focus on using the FLAME model for speech-driven head avatar animation. Recent head animation techniques can be categorized into two main approaches: NeRF-based methods and 3DGS-based methods. NeRF-based approaches (Gafni et al. 2021; Zielonka, Bolkart, and Thies 2023; Hong et al. 2022a; Zheng et al. 2023) leverage neural radiance fields to deform facial movements from a canonical space. HeadNeRF (Hong et al. 2022a) introduces a parametric head model that controls facial shape, expression, and albedo under different lighting conditions. With the emergence of 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), serveral approaches (Qian et al. 2024; Xiang et al. 2024; Xu et al. 2024; Dhamo et al. 2023) have explored its application in head avatar modeling. Gaussian Head Avatar (Xu et al. 2023) initializes Gaussians with a neutral mesh head and incorporates MLPs to deform complex facial expressions.
@@
While these methods produce impressive results in creating 3D head avatars, they overlook precise gaze control and do not generalize well across different subjects. In contrast, GazeGaussian emphasizes precise gaze direction control by decoupling facial animations and gaze movement within a two-stream model. Furthermore, we introduce an expression-guided neural renderer designed to improve the quality of synthesis.
Overview
@@
Figure 2: Pipeline of GazeGaussian. We initialize face-only and eye regions from a pre-trained neutral mesh. Using target expression codes, head rotation, and gaze direction, GazeGaussian optimizes face deformation and eye rotation fields to transform the neutral Gaussians. The transformed Gaussians are splatted into feature maps. The expression codes guide the neural renderer through cross-attention, enabling the rendering of feature maps into high-fidelity images, which are then supervised by multi-view RGB images.
@@
The pipeline of GazeGaussian is illustrated in Fig. 2, including the two-stream Gaussians and the proposed expression-guided neural renderer. Before the beginning of the pipeline, we follow the data preprocessing in GazeNeRF (Ruzzi et al. 2023) and Gaussian Head Avatar (Xu et al. 2024), which include background removal, gaze direction normalization, and facial tracking for each frame. To obtain a neutral mesh for Gaussian initialization, we first reconstruct a Sign Distance Function (SDF) based neutral geometry and then optimize a face deformation field and an eye rotation field from the training data. A neutral mesh representing a coarse geometry across different subjects can be extracted using DMTet (Shen et al. 2021). We then partition the neutral mesh into face-only and eye regions using 3D landmarks, initializing the two-stream Gaussians. Based on these neutral Gaussians, GazeGaussian optimizes a face deformation field and an eye rotation field to transform the Gaussians according to the target expression codes, gaze direction, and head rotation. Next, we concatenate the two-stream Gaussians and rasterize them into a high-dimensional feature map representing the head, face-only, and eye regions. Finally, these feature maps are fed into the expression-guided neural renderer to generate high-fidelity gaze redirection images. The ground truth image is used to supervise the rendered face-only, head, and eye images.
Method
Preliminaries
The vanilla 3D Gaussians (Kerbl et al. 2023) with $N$ points are represented by their positions $X$, the multi-channel color $C$, the rotation $Q$, scale $S$ and opacity $A$. The color $C$ is computed using spherical harmonics, and the rotation $Q$ is represented as the quaternion. These Gaussians are then rasterized and rendered to a multi-channel image $I$ based on the camera parameters $\mu$. This rendering process can be expressed as:
$I=\mathcal{R}(X,C,Q,S,A;\mu),$ (1)
Two-stream GazeGaussian Representation
@@
Our task is to synthesize a head avatar conditioned on gaze direction, head rotation, and expression latent codes. To decouple the complex movements in the face and eyes, we introduce a two-stream Gaussian model consisting of a face-only branch and an eye branch. In the following subsections, we will describe the face deformation and eye rotation processes, respectively.
Face Deformation
For the face-only branch, inspired by Gaussian Head Avatar, we first construct canonical neutral face Gaussians with attributes: $\{\boldsymbol{X}_{0}^{f},\boldsymbol{F}_{0}^{f},\boldsymbol{Q}_{0}^{f},\boldsymbol{S}_{0}^{f},\boldsymbol{A}_{0}^{f}\}$, which are fully optimizable. $\boldsymbol{X}_{0}^{f}\in\mathbb{R}^{N\times 3}$ represents the positions of the Gaussians in the canonical space. $\boldsymbol{F}_{0}^{f}\in\mathbb{R}^{N\times 128}$ denotes the point-wise feature vectors as their intrinsic properties. $\boldsymbol{Q}_{0}^{f}\in\mathbb{R}^{N\times 4}$, $\boldsymbol{S}_{0}^{f}\in\mathbb{R}^{N\times 3}$ and $\boldsymbol{A}_{0}^{f}\in\mathbb{R}^{N\times 1}$ denotes the neutral rotation, scale and opacity respectively. The neutral color is directly predicted from the point-wise feature vectors $\boldsymbol{F}_{0}^{f}$. Then we construct several MLPs, denoted as $\boldsymbol{\Phi}^{f}$, to serve as face deformation fields that transform the neutral face Gaussians. Next, we describe the process of applying offsets to each Gaussian attribute.
Positions $X^{f}$ of the Gaussians. We predict the displacements respectively controlled by the latent codes and the head pose in the canonical space through two different MLPs: ${}_{def}^{exp}\mathcal{F}^{f}\in\boldsymbol{\Phi}^{f}$ and ${}_{def}^{pose}\mathcal{F}^{f}\in\boldsymbol{\Phi}^{f}$. Then, we add them to the neutral positions.
$\begin{split}X^{f}=\boldsymbol{X}_{0}^{f}&+\lambda_{exp}(\boldsymbol{X}_{0}^{f})_{def}^{exp}\mathcal{F}^{f}(\boldsymbol{X}_{0}^{f},\theta)\\ &+\lambda_{pose}(\boldsymbol{X}_{0}^{f})_{def}^{pose}\mathcal{F}^{f}(\boldsymbol{X}_{0}^{f},\beta),\end{split}$ (2)
$\theta$ denoting latent codes including expression and identity coefficients and $\beta$ denoting the head pose. $\lambda_{exp}(\cdot)$ and $\lambda_{pose}(\cdot)$ represent the degree to which the point is influenced by the expression or head pose, respectively, which can be calculated as:
$\begin{split}\lambda_{exp}(x)=\left\{\begin{array}[]{ll}1,&dist(x,\boldsymbol{P}_{0}^{f})<t_{1}\\ \frac{t_{2}-dist(x,\boldsymbol{P}_{0}^{f})}{t_{2}-t_{1}},&dist(x,\boldsymbol{P}_{0}^{f})\in[t_{1},t_{2}]\\ 0,&dist(x,\boldsymbol{P}_{0}^{f})>t_{2}\end{array}\right.\end{split}$
with $\lambda_{pose}(x)=1-\lambda_{exp}(x)$, where $x\in\mathbf{X}_{0}^{f}$ denotes the position of a neutral Gaussian, $dist(x,\boldsymbol{P}_{0}^{f})$ represents the minimum distance from point $x$ to the 3D landmarks (without eyes) $\boldsymbol{P}_{0}^{f}$. Following the approach in Gaussian Head Avatar, the predefined hyperparameters are set as $t_{1}=0.15$ and $t_{2}=0.25$.
Color $C^{f}$ of the Gaussians. Modeling dynamic details requires a color that varies with expressions. The color is directly predict by two color MLPs: ${}_{col}^{exp}\mathcal{F}^{f}\in\boldsymbol{\Phi}^{f}$ and ${}_{col}^{pose}\mathcal{F}^{f}\in\boldsymbol{\Phi}^{f}$:
$\begin{split}C^{f}&=\lambda_{exp}(\boldsymbol{X}_{0}^{f})_{col}^{exp}\mathcal{F}^{f}(\boldsymbol{F}_{0}^{f},\theta)\\ &+\lambda_{pose}(\boldsymbol{X}_{0}^{f})_{col}^{pose}\mathcal{F}^{f}(\boldsymbol{F}_{0}^{f},\beta),\end{split}$ (3)
Rotation, Scale and Opacity $\{Q^{f},S^{f},A^{f}\}$ of the Gaussians. These three attributes are also dynamic, capturing detailed expression-related appearance changes. We just use another two attribute MLPs: ${}_{att}^{exp}\mathcal{F}^{f}\in\boldsymbol{\Phi}^{f}$ and ${}_{att}^{pose}\mathcal{F}^{f}\in\boldsymbol{\Phi}^{f}$ to predict their shift from the neutral value.
$\begin{split}\{Q^{f},S^{f},A^{f}\}&=\{\mathbf{Q}_{0}^{f},\mathbf{S}_{0}^{f},\mathbf{A}_{0}^{f}\}\\ &+\lambda_{exp}(\boldsymbol{X}_{0}^{f})_{att}^{exp}\mathcal{F}^{f}(\boldsymbol{F}_{0}^{f},\theta)\\ &+\lambda_{pose}(\boldsymbol{X}_{0}^{f})_{att}^{pose}\mathcal{F}^{f}(\boldsymbol{F}_{0}^{f},\beta),\end{split}$ (4)

### v2

@@ Abstract
Gaze estimation encounters generalization challenges when dealing with out-of-distribution data. To address this problem, recent methods use neural radiance fields (NeRF) to generate augmented data. However, existing methods based on NeRF are computationally expensive and lack facial details. 3D Gaussian Splatting (3DGS) has become the prevailing representation of neural fields. While 3DGS has been extensively examined in head avatars, it faces challenges with accurate gaze control and generalization across different subjects. In this work, we propose GazeGaussian, the first high-fidelity gaze redirection method that uses a two-stream 3DGS model to represent the face and eye regions separately. Leveraging the unstructured nature of 3DGS, we develop a novel representation of the eye for rigid eye rotation based on the target gaze direction. To enable synthesis generalization across various subjects, we integrate an expression-guided module to inject subject-specific information into the neural renderer. Comprehensive experiments show that GazeGaussian outperforms existing methods in rendering speed, gaze redirection accuracy, and facial synthesis across multiple datasets. The code is available at: https://ucwxb.github.io/GazeGaussian.
@@ 1 Introduction
Figure 1: Gaze redirection: Given an input image, GazeGaussian deforms face and eye Gaussians from canonical space to generate high-fidelity head images with accurate gaze redirection.
@@
Gaze estimation is a fundamental component across various applications [1, 29, 27], yet current estimators [4, 3, 50] often struggle to generalize effectively to out-of-distribution data. To address this, recent approaches [36, 52, 42] have started exploring gaze redirection, which manipulates the gaze in an input image toward a target direction. This process generates augmented data to enhance the generalization capabilities of gaze estimators.
@@
Earlier methods [10, 53, 54, 57] formulate gaze redirection as a 2D image manipulation task, relying on deep learning techniques to warp eye regions of the image toward the target gaze direction. However, these 2D approaches overlook the inherently 3D nature of head and gaze manipulation, often resulting in poor spatial consistency and limited synthesis fidelity. With advancements in Neural Radiance Fields (NeRF) [28] and its variants [44, 46], several methods [62, 12, 60, 16] have achieved 3D dynamic head representation and high-fidelity avatar synthesis. Meanwhile, to enable precise control of gaze direction, recent research [36, 52, 42] has introduced approaches that decouple the face and eye regions, modeling each with separate neural fields to achieve accurate gaze redirection.
@@
As NeRF-based methods are hindered by high computational demands, 3D Gaussian Splatting [18] and its variants [25, 17, 45] achieve impressive rendering quality with significantly faster training speeds. Recent research [49, 33, 51] has applied these methods to 3D head animation, typically using face-tracking [41, 61] parameters to model dynamic 3D head representations. However, existing 3DGS-based approaches neglect the accurate control of gaze direction and struggle to generalize across different subjects, limiting their effectiveness for gaze redirection.
@@
To address the above issues, we propose GazeGaussian, a high-fidelity gaze redirection method that leverages a two-stream 3D Gaussian Splatting (3DGS) model to represent the face and eye regions, respectively. To our knowledge, this is the first integration of 3DGS into gaze redirection tasks. An overview is shown in Fig. 1.
@@
GazeGaussian begins by initializing the two-stream 3DGS model using a pre-trained neutral mesh on the training dataset. This mesh is divided into distinct regions for the face and eyes. By employing gaze direction and face tracking codes, we optimize a deformation field for the face and a rotation field for the eyes, allowing us to adjust the neutral Gaussians accordingly. To achieve precise eye rotation aligned with the target gaze, we present a novel Gaussian Eye Rotation Representation. In contrast to methods like GazeNeRF that implicitly alter feature maps, GazeGaussian explicitly adjusts the position of Gaussians in the eye branch according to the desired gaze direction, utilizing the controllable nature of 3DGS. To address possible errors in gaze direction, GazeGaussian develops an eye rotation field to enhance redirection accuracy. The two-stream Gaussians are rasterized into high-level features and sent to the neural renderer. Finally, to enable the generalization ability across different subjects and inject subject-specific information, we employ an expression-guided neural renderer (EGNR) to synthesize the final gaze-redirection images.
@@
  • We introduce GazeGaussian, the first 3DGS-based gaze redirection pipeline, achieving precise gaze manipulation and generalizable head avatar synthesis.
@@
  • To equip 3DGS with generalization ability, we design an expression-guided neural renderer (EGNR) to integrate subject-specific characteristics.
@@ 2 Related Work
Gaze Redirection. Gaze redirection is the task of manipulating the gaze direction of a face image to a target direction while preserving the subject’s identity and other facial details. Earlier approaches for gaze redirection include novel view synthesis [5, 22, 11], eye-replacement [34, 38], and warping-based methods [10, 20, 47]. However, these methods are limited by person-specific data, restricted redirection range, and artifact introduction. To further improve gaze redirection, recent studies [30, 59, 14, 48] have employed neural network-based generative models. STED [59], building on FAZE [30], introduces a self-transforming encoder-decoder that generates full-face gaze redirection images. With advancements in Neural Radiance Fields (NeRF) [28], several studies [23, 36, 52, 42] have aimed to model the complex rotation of the eyeball. GazeNeRF [36] employs a two-stream MLP architecture to separately model the face and eye regions, achieving improved gaze redirection performance.
@@
However, these methods are hindered by substantial computational demands and limited rendering efficiency. Gaze manipulation occurs at the feature maps and remains an implicit approach. In contrast, GazeGaussian allows for explicit control over eye rotations, improving gaze redirection accuracy and accelerating the synthesis process.
@@
Head Avatar Synthesis. The synthesis of head avatars has garnered considerable attention in recent years. FLAME [24] is a parameterized 3D head model that maps parameters of shape, expression, and pose onto 3D facial geometry, allowing for controllable head avatar generation. Many subsequent works [6, 35, 8, 31, 32, 2] focus on using the FLAME model for speech-driven head avatar animation. Recent head animation techniques can be categorized into two main approaches: NeRF-based methods and 3DGS-based methods. NeRF-based approaches [9, 62, 16, 60] leverage neural radiance fields to deform facial movements from a canonical space. HeadNeRF [16] introduces a parametric head model that controls facial shape, expression, and albedo under different lighting conditions. With the emergence of 3D Gaussian Splatting (3DGS) [18], several approaches [33, 49, 51, 7] have explored its application in single head avatar modeling. Gaussian Head Avatar [51] initializes Gaussians with a neutral mesh head and incorporates MLPs to deform complex facial expressions.
@@
However, these methods ignore precise gaze control and are confined to the single avatar animation. In contrast, GazeGaussian achieves precise gaze direction control by decoupling facial animations and gaze movement within a two-stream model. Furthermore, an expression-guided neural renderer is designed to inject person-specific information for generalization ability across different subjects.
3 Overview
@@
Figure 2: Pipeline of GazeGaussian. We initialize face-only and eye regions from a pre-trained neutral mesh. Using target expression codes, head rotation, and gaze direction, GazeGaussian optimizes face deformation and eye rotation fields to transform the neutral Gaussians. The transformed Gaussians are splatted into feature maps. The expression codes guide the neural renderer through cross-attention, enabling the rendering of feature maps into high-fidelity gaze redirection images, which are then supervised by multi-view ground truths.
@@
The pipeline of GazeGaussian is illustrated in Fig. 2, including the two-stream Gaussians and the proposed expression-guided neural renderer. Before the beginning of the pipeline, we follow the data preprocessing in GazeNeRF [36] and Gaussian Head Avatar [51], which include background removal, gaze direction normalization, and facial tracking for each frame. To obtain a neutral mesh for Gaussian initialization, we first reconstruct a Sign Distance Function (SDF) based neutral geometry and then optimize a face deformation field and an eye rotation field from the training data. A neutral mesh representing a coarse geometry across different subjects can be extracted using DMTet [37]. We then partition the neutral mesh into face-only and eye regions using 3D landmarks, initializing the two-stream Gaussians. Based on these neutral Gaussians, GazeGaussian optimizes a face deformation field and an eye rotation field to transform the Gaussians according to the target expression codes, gaze direction, and head rotation. Next, we concatenate the two-stream Gaussians and rasterize them into a high-dimensional feature map representing the head, face-only, and eye regions. Finally, these feature maps are fed into the expression-guided neural renderer to generate high-fidelity gaze redirection images. The ground truth image is used to supervise the rendered face-only, head, and eye images.
4 Method
4.1 Preliminaries
The vanilla 3D Gaussians [18] with $n$ points are represented by their center $\mu\in\mathbb{R}^{3}$, the color computed by $k$ degrees of spherical harmonics $c\in\mathbb{R}^{3(k+1)^{2}}$, the quaternion rotation $R\in\mathbb{R}^{4}$, scale $S\in\mathbb{R}^{3}$ and opacity $\alpha\in\mathbb{R}$. The covariance matrix $\Sigma$ can be obtained by $\Sigma=RSS^{T}R^{T}$, which then formulates the Gaussian function $G(x)=e^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)}$. These Gaussians are then rasterized into an image $I$, in which the pixel color $C$ is computed by blending $N$ ordered Gaussians overlapping the pixel:
$C=\sum^{N}_{i=1}{c_{i}\alpha_{i}\prod_{j=1}^{i-1}}(1-\alpha_{j}),$ (1)
4.2 Two-stream GazeGaussian Representation
@@
Our task is to synthesize a head avatar conditioned on gaze direction, head rotation, and expression latent codes. To decouple the complex movements in the face and eyes, we introduce a two-stream Gaussian model consisting of a face-only deformation branch and an eye rotation branch in the following section.
4.2.1 Face Deformation
For the face-only branch, inspired by Gaussian Head Avatar [51], we first construct canonical neutral face Gaussians with trainable attributes: $\{\mu_{0}^{f},z_{0}^{f},R_{0}^{f},S_{0}^{f},\alpha_{0}^{f}\}$. All attributes are identical to a standard Gaussian, except for $z_{0}^{f}\in\mathbb{R}^{128}$, which signifies the point-wise feature vectors. Rather than directly modeling complex color $c^{f}$, we decode $z_{0}^{f}$ into $c^{f}$ and affect other attributes’ deformation to capture intricate facial details.
To transform canonical face Gaussians into the target space, we first model the influence of latent codes $\tau$ and head pose $\gamma$ through respective weights $\lambda_{\tau}$ and $\lambda_{\gamma}$. We compute the minimum distance $d$ from each Gaussian center $\mu$ to the 3D facial landmarks (excluding eye regions). Based on the distance, we assign influence weights $\lambda_{\tau}$ and $\lambda_{\gamma}$ to the expressive latent codes $\tau$ and head pose $\gamma$. For Gaussians near landmarks ($d<d_{1}$), they are primarily influenced by expressive latent codes and set $\lambda_{\tau}=1$. For distant Gaussians ($d>d_{2}$), they are predominantly affected by head pose, with $\lambda_{\tau}=0$. To ensure a smooth transition between these regions ($d_{2}\geq d\geq d_{1}$), we define $\lambda_{\tau}=(d_{2}-d)/(d_{2}-d_{1})$. Consequently, the pose weight is determined as $\lambda_{\gamma}=1-\lambda_{\tau}$. The distance thresholds are empirically set to $d_{1}=0.15$ and $d_{2}=0.25$.
After that, we develop MLPs $\{E_{\mu}^{f},E_{c}^{f},E_{\kappa}^{f}\}\in E^{f}$ for latent codes-conditioned expression and $\{P_{\mu}^{f},P_{c}^{f},P_{\kappa}^{f}\}\in P^{f}$ for head pose-conditioned deformation:
$\begin{split}\mu^{f}&=\mu_{0}^{f}+\lambda_{\tau}E_{\mu}^{f}(\mu_{0}^{f},\tau)+\lambda_{\gamma}P_{\mu}^{f}(\mu_{0}^{f},\gamma),\\ c^{f}&=\lambda_{\tau}E_{c}^{f}(z_{0}^{f},\tau)+\lambda_{\gamma}P_{c}^{f}(z_{0}^{f},\gamma),\\ \kappa^{f}&=\kappa_{0}^{f}+\lambda_{\tau}E_{\kappa}^{f}(z_{0}^{f},\tau)+\lambda_{\gamma}P_{\kappa}^{f}(z_{0}^{f},\gamma),\end{split}$ (2)